• A certain substance can cool itself from 110°C to 70°C in 15minutes. If the air’s temperature is 30°C, how long will it take to cool from 110°C to 50°C?
1
Expert's answer
2021-07-09T07:24:37-0400

Newton's Law of Cooling 


"\dfrac{dT}{dt}=k(T-T_s)"

"\dfrac{dT}{T-T_s}=kdt"

"T=T_s+Ce^{kt}"



"T_s=30\degree C, T(0)=110\degree C"



"110=30+C=>C=80\degree C"


"T=30+80e^{kt}"



"T(0.25)=70\degree C"


"70=30+80e^{0.25k}"

"e^{0.25k}=\dfrac{1}{2}"

"k=-4\ln 2"




"T=30+80e^{(-4\ln2) t}"

If "T=50\degree C"


"50=30+80e^{(-4\ln2) t}"

"e^{(-4\ln2) t}=\dfrac{1}{4}"

"(-4\ln2) t=-2\ln2"

"t=0.5\ hr"

30 minutes.